A steel rod 4 m long and 0.5 cm2





in cross sectional area is found to stretch 0.2





cm under a tension of 12000 N. Under these conditions, determine




(a) The stress




(b) The strain; and




(c) Young’s modulus for the material

1
Expert's answer
2022-03-01T11:12:23-0500

Given:

"l_0=4\:\rm m"

"A=0.5*10^{-4}\:\rm m^2"

"\Delta l=0.002\:\rm m"

"F=12000\:\rm N"

(a) The stress

"\sigma=\frac{F}{A}=\frac{12000}{0.5*10^{-4}}=2.4*10^8\:\rm Pa"

(b) The strain

"\varepsilon=\frac{\Delta l}{l_0}=\frac{0.002}{4}=5*10^{-4}"

(c) The Young’s modulus for the material

"E=\frac{\sigma}{\varepsilon}=\frac{2.4*10^8}{5*10^{-4}}=4.8*10^{11}\:\rm Pa"